OpenStudy (anonymous):

Find the roots of the polynomial equation. x2 + 2x – 48 = 0 A. –6, 8 B. –4, 12 C. –12, –4 D. –8, 6 @elissaab

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b^2-4*a*c} }{ 2a }\]

OpenStudy (anonymous):

does that look familiar to you?

OpenStudy (mathmale):

x2 + 2x – 48 = 0, or \[ x^2 + 2x – 48 = 0\] is a quadratic equation. Each of the three terms on the left has a constant coefficient. What are the three coefficients, a, b and c? Once you've identified those, plug them into the quadratic formula, as given you by zimmah, above, and determine the values of the roots of this quadratic equation.

OpenStudy (anonymous):

In this case by the way there is an easier method, but it's not always as easy. you can try to find two numbers so that: 1) when you add up these two numbers they make +2 2) when you multiply these two numbers they make -48

OpenStudy (anonymous):

Since you tagged me, here is a hint http://www.wolframalpha.com/input/?i=factor++x%5E2+%2B+2x+%E2%80%93+48